# Opinion Dynamics for Utility Maximizing Agents: Exploring the Impact of Resource Penalty

### v1

@@ 5.2 Nash Equilibria
Here we carry out a Nash Equilibrium analysis of the opinion formation game. Recall that every agent is interested in maximizing its utility $U_{i}$ given in (1) by suitably choosing its opinion $z_{i}$. Thus, this gives a strategic form game $\mathcal{G}=\langle\mathcal{V},\left(\mathbb{R}\right)_{i\in\mathcal{V}},\left(U_{i}\right)_{i\in\mathcal{V}}\rangle$ among the set of agents $\mathcal{V}$, with the strategy of agent $i\in\mathcal{V}$ being its opinion $z_{i}\in\mathbb{R}$ and its utility function being $U_{i}(\cdot)$. For the sake of convenience, we let $\mathbf{z}_{-i}$ denote the opinions of all agents other than $i$. Then, the set of Nash equilibria of the game $\mathcal{G}$ is
$\mathcal{NE}(\mathbf{p})$ :=$:=\{\mathbf{z}^{*}\in\mathbb{R}^{n}\,\,|\,\,\forall i\in\mathcal{V},$
$\quad U_{i}(z^{*}_{i},\mathbf{z}_{-i}^{*},p_{i})\geq U_{i}(z_{i},\mathbf{z}_{-i}^{*},p_{i}),\forall z_{i}\in\mathbb{R}\}\,.$ (14)
Note that for a Nash equilibrium $\mathbf{z}^{*}$, $z^{*}_{i}$ is agent $i$’s best response over all opinions $z_{i}\in\mathbb{R}$ to $\mathbf{z}_{-i}^{*}$, the opinion profile of all the other agents. However, in the dynamics (2), each agent updates its opinion according to the gradient ascent of its utility with respect to its opinion while assuming that the other agents do not change their opinions. Hence, the agents at each time instant revise their opinion only towards the “local” best response. This motivates the next definition.
@@
(Local Nash equilibrium). A strategy profile $\mathbf{z}^{*}\in\mathbb{R}^{n}$ is said to be a local Nash equilibrium if and only if $\forall i\in\mathcal{V}$, $\exists\,\,\rho_{i}\in\mathbb{R}_{>0}$ such that
@@
The set of local Nash equilibria of $\mathcal{G}$ is denoted by $\mathcal{NE}_{l}(\mathbf{p})$. normal-$\bullet$
@@
(Local Nash equilibrium is an equilibrium of the opinion dynamics). If an opinion profile $\mathbf{z}^{*}$ is such that $\mathbf{z}^{*}\in\mathcal{NE}_{l}$ then $\mathbf{z}^{*}\in\mathcal{E}$.
@@
Lemma 5.8 states that every local Nash equilibrium of $\mathcal{G}$ is also an equilibrium point of dynamics (2). But the converse need not be true. In the following result, we give conditions for an opinion profile $\mathbf{z}^{*}\in\mathbb{R}^{n}$ that is an equilibrium point of the opinion dynamics model to be a local Nash equilibrium of the opinion formation game.
@@
(Conditions for an equilibrium point of (2) to be a local Nash equilibrium). Consider the dynamics (2) and the set of equilibrium points $\mathcal{E}$ in (8). Let $\mathbf{z}^{*}=\left(z_{i}^{*},\mathbf{z}_{-i}^{*}\right)\in\mathcal{E}$. Then, $\mathbf{z}^{*}\in\mathcal{NE}_{l}$ only if for each $i\in\mathcal{V}$,
$(z_{i}^{*})^{2}\geq\tau_{i}:=\frac{r_{i}}{3}\left[\sum_{k\in\mathcal{N}_{i}^{e}}\frac{|a_{ik}|r_{k}}{B}-\sum_{k\in\mathcal{N}_{i}^{f}}\frac{a_{ik}r_{k}}{B}-\frac{w_{i}r_{i}}{B}\right].$ (15)
@@
Let the hypothesis be true. From Definition 5.7 we know that $\mathbf{z}^{*}=\left(z_{i}^{*},\mathbf{z}_{-i}^{*}\right)$ is a local Nash equilibrium if and only if $\forall i\in\mathcal{V}$, $z_{i}^{*}$ locally maximizes $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$. Now since $\mathbf{z}^{*}\in\mathcal{E}$, by the definitions in (2) and (8), it is clear that for each $i\in\mathcal{V}$, $z_{i}^{*}$ satisfies the first-order necessary conditions for it to be a local maximizer of $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$. Now, suppose that $\mathbf{z}^{*}\in\mathcal{NE}_{l}$. Then, we have that
@@
(Non-existence of local Nash equilibria). Suppose $\eta_{i}>0$ is an ultimate bound on $z_{i}$ for any $i\in\mathcal{V}$ under (2). If there exists $i\in\mathcal{V}$ such that $\tau_{i}>\eta_{i}^{2}$, with $\tau_{i}$ defined in (15), then $\mathcal{NE}_{l}=\varnothing$. normal-$\blacksquare$
@@
(Equality of equilibria sets and uniqueness of Nash equilibrium). Consider the dynamics in (3a) and the set of equilibrium points $\mathcal{E}$ in (8). Suppose for each agent $i\in\mathcal{V}$, $\tau_{i}\leq 0$. Then, $\mathcal{NE}=\mathcal{NE}_{l}=\mathcal{E}$. Moreover, if Assumption (A4) holds then it is a singleton set.
@@
Suppose that $\tau_{i}\leq 0$, $\forall i\in\mathcal{V}$. First, it is obvious that $\mathcal{NE}\subseteq\mathcal{NE}_{l}$ and Lemma 5.8 implies that $\mathcal{NE}_{l}\subseteq\mathcal{E}$. Next, we show that $\mathcal{E}\subseteq\mathcal{NE}_{l}$. Consider an equilibrium $\mathbf{z}^{*}\in\mathcal{E}$. It can be easily verified that for each $i\in\mathcal{V}$, $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is strictly concave in $z_{i}$. Since $\mathbf{z}\in\mathcal{E}$, the first order necessary condition for a point to be a local maximizer of $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is satisfied for each $i$. Hence, $\mathbf{z}^{*}\in\mathcal{NE}_{l}$. Finally, we show that $\mathcal{NE}_{l}\subseteq\mathcal{NE}$. Since $\forall i\in\mathcal{V}$, $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is a strictly concave function for each $\mathbf{z}_{-i}^{*}$, the implication follows directly from the definitions of $\mathcal{NE}$ and $\mathcal{NE}_{l}$. This completes the proof of the first part of the claim. Now suppose that Assumption (A4) holds. Then again $\tau_{i}<0$, $\forall i\in\mathcal{V}$. Uniqueness of Nash equilibrium now follows from Theorem 4.5 and first part of the claim.
@@
Recall the interpretation of $w_{i}$ as the weight of influence of a fully stubborn virtual agent on agent $i$ provided in Remark 3.2. Using this interpretation, let us define the weighted in-degree for each agent $j\in\mathcal{V}$ in social network $\mathsf{G}$ as $d_{j}^{in}:=\left[\frac{w_{j}r_{j}}{B}+\sum_{k\in\mathcal{V}\setminus\left\{j\right\}}\frac{a_{jk}r_{k}}{B}\right]$. Then for each $j\in\mathcal{V}$, $\tau_{j}=-r_{j}d_{j}^{in}/3$ and $\tau_{j}<0(=0)$ if and only if $d_{j}^{in}>0(=0)$. The condition $d_{j}^{in}>0$, $\forall j\in\mathcal{V}$ can be interpreted as follows: For each agent $i\in\mathcal{V}$, the aggregate influence weight of its friends (including the virtual fully stubborn agent since $w_{i}>0$) is greater than the aggregate influence weight of its enemies. If this holds true then Theorem 5.13 states that each equilibrium point of the dynamics is also a Nash equilibrium of the game. Next, if Assumption (A4) holds then for each $j\in\mathcal{V}$, $d_{j}^{in}>0$ and hence $\tau_{j}<0$. In other words, the opinion formation game has a unique Nash equilibrium if the willingness of each agent is more than twice the aggregate influence weight of its enemies. Moreover, thanks to Theorem 4.5, we can say that the opinions under (2) always converge to this unique Nash equilibrium, starting from any initial condition. normal-$\bullet$
@@ 5.3 Price of Anarchy
$\mathsf{C}_{E}(\mathbf{z},\mathbf{p})$ $:=\max_{i\in\mathcal{V}}\chi_{i}(\mathbf{z},\mathbf{p})=\max_{i\in\mathcal{V}}-U_{i}(\mathbf{z},\mathbf{p}),$ (17)
$\mathsf{C}_{U}(\mathbf{z},\mathbf{p})$ $:=\sum_{i\in\mathcal{V}}\chi_{i}(\mathbf{z},\mathbf{p})=-\sum_{i\in\mathcal{V}}U_{i}(\mathbf{z},\mathbf{p}),$ (18)
@@
(Price of anarchy). Consider the opinion formation game $\mathcal{G}$ corresponding to the opinion dynamics (2), and $\mathcal{NE}$, the set of its Nash equilibria. Suppose Assumption (A3) holds and $\mathbf{p}\neq 0$. The egalitarian and the utilitarian prices of anarchy, $\pi_{e}$ and $\pi_{u}$, respectively are defined as
@@
    (Non-zero preferences). $\forall i\in\mathcal{V}$, $p_{i}\neq 0$. $\bullet$
@@
(Satisfaction ratio). Consider the cost function $\chi_{i}(\mathbf{z})$ of each agent $i\in\mathcal{V}$. Suppose Assumptions (A3) and (A5) hold. The satisfaction ratio $\mathsf{SR}_{i}$ for any agent $i\in\mathcal{V}$ at any $\mathbf{z}\in\mathbb{R}^{n}$ is defined as,
@@
(Convexity of cost function). Suppose that Assumption (A3) holds. Then for each $i\in\mathcal{V}$, $\chi_{i}(\cdot)$ is convex with $m_{i}\mathbf{1}$ as one of its minimizer. normal-$\bullet$
@@
(Bounds on prices of anarchy). Consider the dynamics (2). Suppose Assumptions (A3) and (A5) hold. Let us denote the unique Nash equilibrium of the game underlying the opinion dynamics (2) by $\mathbf{z}^{*}\in\mathcal{NE}$. Then,
@@
with $\pi_{e}$ and $\pi_{u}$ as defined in Definition 5.16. normal-$\bullet$
@@
($\mathsf{PoA}$ is unity for a non-neutral consensus equilibrium). Consider the dynamics (2). Suppose Assumptions (A3) and (A5) hold. Suppose the unique $\mathbf{z}^{*}\in\mathcal{NE}=\mathcal{E}$ is a non-neutral consensus equilibrium, i.e., $\mathbf{z}^{*}=m\mathbf{1}$ for some $m\neq 0$. Then, $\pi_{e}=\pi_{u}=1$.
@@
(Neutral consensus is socially optimal). Suppose Assumption (A3) holds and $\mathbf{p}=\mathbf{0}$. From Theorems 5.14.5 and 5.13, we know that the unique equilibrium point $\mathbf{z}^{*}=\mathbf{0}\in\mathcal{E}=\mathcal{NE}$. Now, from Lemma 5.18 we know that $\mathbf{0}$ is a minimizer of $\chi_{i}(\mathbf{z})$, $\mathsf{C}_{E}(\mathbf{z})$ and $\mathsf{C}_{U}(\mathbf{z})$. Thus prices of anarchy (19) and satisfaction ratios (20) are not defined in this case. But since $\mathsf{C}_{E}(\mathbf{z}^{*})=\min_{\mathbf{z}\in\mathbb{R}^{n}}\mathsf{C}_{E}(\mathbf{z})=0=\mathsf{C}_{U}(\mathbf{z}^{*})=\min_{\mathbf{z}\in\mathbb{R}^{n}}\mathsf{C}_{U}(\mathbf{z})$, neutral consensus equilibrium is also socially optimal. normal-$\bullet$
@@ 6 Oscillatory Behavior of Opinions
In [26], we studied a special case of the dynamics proposed in the current paper. We showed through numerical simulations that in certain scenarios, the opinions exhibit oscillatory behavior. In this section, we analyze such behavior in the case of a pair of agents $\mathcal{V}=\{1,2\}$. For a general social network with $n$ agents, the analysis is significantly more complicated due to the higher dimension of the state space but also due to the complexity added by the social network. Thus, the analysis in the general case of a social network with $n$ agents would have to be a separate research work in itself and is out of the scope of this paper.
@@
$\begin{split}\dot{z}_{1}&=S_{1}(z_{1})+\big{[}c_{1}r_{2}/{B}\big{]}\big{[}z_{2}-z_{1}\big{]}\\ \dot{z}_{2}&=S_{2}(z_{2})-\big{[}c_{2}r_{1}/{B}\big{]}\big{[}z_{2}-z_{1}\big{]}.\end{split}$ (22)
We call an agent $i\in\mathcal{V}$ a contrarian (non-contrarian) if its conformity weight $c_{i}<0\ (\geq 0)$. For the two-agent dynamics (22), it can be easily verified that if the opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit oscillatory behavior then they will have the same fundamental period of oscillations. We state this in the following result (proof in Appendix).
@@
(Equal period of oscillations) Consider the two-agent dynamics (22). Suppose the opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit oscillations with fundamental periods $T_{1}>0$ and $T_{2}>0$, respectively. Then, $T_{1}=T_{2}$. normal-$\bullet$
Note that periodic orbits can exist for (22) only if at least one of the agents is a contrarian as otherwise Theorem 4.5 guarantees convergence of opinions to the unique equilibrium from any arbitrary initial condition. In the next result (proof in Appendix), we give a necessary condition for the existence of periodic orbits for (22).
@@
(Necessary condition for existence of periodic solutions) Consider the two-agent dynamics (22). Opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit periodic behavior only if all the following hold: (i) $c_{1}<0$ or $c_{2}<0$, (ii) $c_{1}c_{2}\neq 0$, and (iii) $(c_{2}r_{1}+w_{1}r_{1}+w_{2}r_{2}+c_{1}r_{2})<0$. normal-$\bullet$
@@
(Geometric and other interpretations of the conditions for the existence of periodic opinions). The condition given in Lemma 6.2 ensures existence of the following ellipse in the phase plane,
$\frac{z_{1}^{2}}{r_{1}}+\frac{z_{2}^{2}}{r_{2}}=-\frac{(w_{1}r_{1}+w_{2}r_{2}+c_{1}r_{2}+c_{2}r_{1})}{3B}=:\upsilon>0.$ (23)
@@
The oscillatory behavior of opinions appear only when at least one of the agents is a contrarian. This is similar to the so-called boomerang effect [29], where at least one of the agents has an antagonist influence on another agent. As a result, the agent on whom there is an antagonist influence shifts its opinion away from the other agent. In such situations, the opinions of the agents could converge to a disagreement equilibrium or could possibly keep on oscillating forever never converging to an equilibrium. normal-$\bullet$
Now that we have dealt with necessary conditions for a periodic solution of (22) to exist, we conclude this section by giving sufficient conditions for the agents to exhibit periodic opinion profiles. More specifically, we give sufficient conditions on the conformity weights $c_{1}$, $c_{2}$ for a Hopf bifurcation to exist (proof in Appendix).
@@
(Existence of Hopf bifurcation) Consider two-agent dynamics (22) with $c_{2}\in\mathbb{R}$ as the bifurcation parameter. Let $g:=r_{1}/B>0$, $d:=(c_{1}r_{2})/B\in\mathbb{R}$ and $\kappa_{i}:=\left[\frac{w_{i}r_{i}}{B}+\frac{3m_{i}^{2}}{r_{i}}\right]>0,i\in\{1,2\}$. Suppose $m_{1}=m_{2}=m\in\mathbb{R}$ and $\left[d(\kappa_{2}-\kappa_{1})-\kappa_{1}^{2}\right]>0$. Then there is a family of periodic orbits of (22) that bifurcates from the consensus equilibrium $\mathbf{z}^{*}=m\mathbf{1}$ at $c_{2}=c_{2}^{*}:=-(\kappa_{1}+\kappa_{2}+d)/g$. normal-$\bullet$

### v2

@@ 5.2 Nash Equilibria
Here we carry out a Nash Equilibrium analysis of the opinion formation game. Recall that every agent is interested in maximizing its utility $U_{i}$ given in (1) by suitably choosing its opinion $z_{i}$. Thus, this gives a strategic form game $\mathcal{G}=\langle\mathcal{V},\left(\right)_{i\in\mathcal{V}},\left(U_{i}\right)_{i\in\mathcal{V}}\rangle$ among the set of agents $\mathcal{V}$, with the strategy of agent $i\in\mathcal{V}$ being its opinion $z_{i}\in$ and its utility function being $U_{i}(\cdot)$. For the sake of convenience, we let $\mathbf{z}_{-i}$ denote the opinions of all agents other than $i$. Then, the set of Nash equilibria of the game $\mathcal{G}$ is
$\mathcal{NE}(\mathbf{p})$ :=$:=\{\mathbf{z}^{*}\in^{n}\,\,|\,\,\forall i\in\mathcal{V},$
$\quad U_{i}(z^{*}_{i},\mathbf{z}_{-i}^{*},p_{i})\geq U_{i}(z_{i},\mathbf{z}_{-i}^{*},p_{i}),\forall z_{i}\in\}\,.$ (11)
Note that for a Nash equilibrium $\mathbf{z}^{*}$, $z^{*}_{i}$ is agent $i$’s best response over all opinions $z_{i}\in$ to $\mathbf{z}_{-i}^{*}$, the opinion profile of all the other agents. However, in the dynamics (2), each agent updates its opinion according to the gradient ascent of its utility with respect to its opinion while assuming that the other agents do not change their opinions. Hence, the agents at each time instant revise their opinion only towards the “local” best response. This motivates the next definition.
@@
(Local Nash equilibrium.) A strategy profile $\mathbf{z}^{*}\in^{n}$ is said to be a local Nash equilibrium if and only if $\forall i\in\mathcal{V}$, $\exists\,\,\rho_{i}\in_{>0}$ such that
@@
The set of local Nash equilibria of $\mathcal{G}$ is denoted by $\mathcal{NE}_{l}(\mathbf{p})$. $\bullet$
@@
(Local Nash equilibrium is an equilibrium of the opinion dynamics.) If an opinion profile $\mathbf{z}^{*}$ is such that $\mathbf{z}^{*}\in\mathcal{NE}_{l}$ then $\mathbf{z}^{*}\in\mathcal{E}$.
@@
Lemma 5.8 states that every local Nash equilibrium of $\mathcal{G}$ is also an equilibrium point of dynamics (2). But the converse need not be true. In the following result, we give conditions for an opinion profile $\mathbf{z}^{*}\in\mathbb{R}^{n}$ that is an equilibrium point of the opinion dynamics to be a local Nash equilibrium of the opinion formation game.
@@
(Conditions for an equilibrium point of (2) to be a local Nash equilibrium.) Consider the dynamics (2) and the set of equilibrium points $\mathcal{E}$ in (8). Let $\mathbf{z}^{*}=\left(z_{i}^{*},\mathbf{z}_{-i}^{*}\right)\in\mathcal{E}$. Then, $\mathbf{z}^{*}\in\mathcal{NE}_{l}$ only if for each $i\in\mathcal{V}$,
$(z_{i}^{*})^{2}\geq\tau_{i}:=\frac{r_{i}}{3}\left[\sum_{k\in\mathcal{N}_{i}^{e}}|a_{ik}|-\sum_{k\in\mathcal{N}_{i}^{f}}a_{ik}-w_{i}\right].$ (12)
@@
Let the hypothesis be true. From Definition 5.7 we know that $\mathbf{z}^{*}=\left(z_{i}^{*},\mathbf{z}_{-i}^{*}\right)$ is a local Nash equilibrium if and only if $\forall i\in\mathcal{V}$, $z_{i}^{*}$ locally maximizes $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$. Now since $\mathbf{z}^{*}\in\mathcal{E}$, by the definitions in (2) and (8), it is clear that for each $i\in\mathcal{V}$, $z_{i}^{*}$ satisfies the first-order necessary conditions for it to be a local maximizer of $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$.
Now, suppose that $\mathbf{z}^{*}\in\mathcal{NE}_{l}$. Then, we have that
@@
(Non-existence of local Nash equilibria.) Suppose $\eta_{i}>0$ is an ultimate bound on $z_{i}$ for any $i\in\mathcal{V}$ under (2). If there exists $i\in\mathcal{V}$ such that $\tau_{i}>\eta_{i}^{2}$, with $\tau_{i}$ defined in (12), then $\mathcal{NE}_{l}=\varnothing$. $\blacksquare$
@@
(Equality of equilibria sets and uniqueness of Nash equilibrium.) Consider the dynamics in (3a) and the set of equilibrium points $\mathcal{E}$ in (8). Suppose for each agent $i\in\mathcal{V}$, $\tau_{i}\leq 0$. Then, $\mathcal{NE}=\mathcal{NE}_{l}=\mathcal{E}$. Moreover, if Assumption (A4) holds then it is a singleton set.
@@
Suppose that $\tau_{i}\leq 0$, $\forall i\in\mathcal{V}$. First, it is obvious that $\mathcal{NE}\subseteq\mathcal{NE}_{l}$ and Lemma 5.8 implies that $\mathcal{NE}_{l}\subseteq\mathcal{E}$. Next, we show that $\mathcal{E}\subseteq\mathcal{NE}_{l}$. Consider an equilibrium $\mathbf{z}^{*}\in\mathcal{E}$. It can be easily verified that for each $i\in\mathcal{V}$, $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is strictly concave in $z_{i}$. Since $\mathbf{z}^{*}\in\mathcal{E}$, the first order necessary condition for a point to be a local maximizer of $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is satisfied for each $i$. Hence, $\mathbf{z}^{*}\in\mathcal{NE}_{l}$. Finally, we show that $\mathcal{NE}_{l}\subseteq\mathcal{NE}$. Since $\forall i\in\mathcal{V}$, $U_{i}(\cdot,\mathbf{z}_{-i}^{*})$ is a strictly concave function for each $\mathbf{z}_{-i}^{*}$, the implication follows directly from the definitions of $\mathcal{NE}$ and $\mathcal{NE}_{l}$. This completes the proof of the first part of the claim. Now suppose that Assumption (A4) holds. Then again $\tau_{i}<0$, $\forall i\in\mathcal{V}$. Uniqueness of Nash equilibrium now follows from Theorem 4.3 and first part of the claim.
@@
We can interpret $w_{i}$ as the weight of influence of a stubborn virtual agent on agent $i$. Using this interpretation, let us define the weighted in-degree for each agent $j\in\mathcal{V}$ in social network $\mathsf{G}$ as $d_{j}^{in}:=\left[w_{j}+\sum_{k\in\mathcal{V}\setminus\left\{j\right\}}a_{jk}\right]$. Then for each $j\in\mathcal{V}$, $\tau_{j}=-r_{j}d_{j}^{in}/3$ and $\tau_{j}<0(=0)$ if and only if $d_{j}^{in}>0(=0)$. The condition $d_{j}^{in}>0$, $\forall j\in\mathcal{V}$ can be interpreted as follows: For each agent $i\in\mathcal{V}$, the aggregate influence weight of its friends (including the virtual fully stubborn agent since $w_{i}>0$) is greater than the aggregate influence weight of its enemies. If this holds true then Theorem 5.13 states that each equilibrium point of the dynamics is also a Nash equilibrium of the game. Next, if Assumption (A4) holds then for each $j\in\mathcal{V}$, $d_{j}^{in}>0$ and hence $\tau_{j}<0$. In other words, the opinion formation game has a unique Nash equilibrium if the willingness of each agent to be close to its internal preference (or stubbornness) is more than twice the aggregate influence weight of its enemies. Moreover, thanks to Theorem 4.3, we can say that the opinions under (2) always converge to this unique Nash equilibrium, starting from any initial condition. $\bullet$
@@ 5.3 Price of Anarchy
$\mathsf{C}_{E}(\mathbf{z},\mathbf{p}):=\max_{i\in\mathcal{V}}\chi_{i}(\mathbf{z},\mathbf{p}),\quad\mathsf{C}_{U}(\mathbf{z},\mathbf{p}):=\sum_{i\in\mathcal{V}}\chi_{i}(\mathbf{z},\mathbf{p}),$
@@
(Price of anarchy.) Consider the opinion formation game $\mathcal{G}$ corresponding to the opinion dynamics (2), and $\mathcal{NE}$, the set of its Nash equilibria. Suppose Assumption (A3) holds and $\mathbf{p}\neq\mathbf{0}$. The egalitarian and the utilitarian prices of anarchy, $\pi_{e}$ and $\pi_{u}$, respectively are defined as
@@
    (Non-zero preferences.) $\forall i\in\mathcal{V}$, $p_{i}\neq 0$. $\bullet$
@@
(Satisfaction ratio.) Consider the cost function $\chi_{i}(\mathbf{z})$ of each agent $i\in\mathcal{V}$. Suppose Assumptions (A3) and (A5) hold. The satisfaction ratio $\mathsf{SR}_{i}$ for any agent $i\in\mathcal{V}$ at any $\mathbf{z}\in\mathbb{R}^{n}$ is defined as,
@@
(Convexity of cost function.) Suppose that Assumption (A3) holds. Then for each $i\in\mathcal{V}$, $\chi_{i}(\cdot)$ is convex with $m_{i}\mathbf{1}$ as one of its minimizer. $\bullet$
@@
(Bounds on prices of anarchy.) Consider the dynamics (2). Suppose Assumptions (A3) and (A5) hold. Let us denote the unique Nash equilibrium of the game underlying the opinion dynamics (2) by $\mathbf{z}^{*}\in\mathcal{NE}$. Then,
@@
with $\pi_{e}$ and $\pi_{u}$ as defined in Definition 5.16. $\bullet$
@@
($\mathsf{PoA}$ is unity for a non-neutral consensus equilibrium.) Consider the dynamics (2). Suppose Assumptions (A3) and (A5) hold. Suppose the unique $\mathbf{z}^{*}\in\mathcal{NE}=\mathcal{E}$ is a non-neutral consensus equilibrium, i.e., $\mathbf{z}^{*}=m\mathbf{1}$ for some $m\neq 0$. Then, $\pi_{e}=\pi_{u}=1$.
@@
(Neutral consensus is socially optimal.) Let Assumption (A3) hold and $\mathbf{p}=\mathbf{0}$. Theorems 5.14.3 and 5.13 imply that the unique equilibrium point $\mathbf{z}^{*}=\mathbf{0}\in\mathcal{E}=\mathcal{NE}$. Now, from Lemma 5.18 we know that $\mathbf{0}$ is a minimizer of $\chi_{i}(\mathbf{z})$, $\mathsf{C}_{E}(\mathbf{z})$ and $\mathsf{C}_{U}(\mathbf{z})$. Thus, prices of anarchy (14) and satisfaction ratios (15) are not defined in this case. But since $\mathsf{C}_{E}(\mathbf{z}^{*})=\min_{\mathbf{z}\in\mathbb{R}^{n}}\mathsf{C}_{E}(\mathbf{z})=0=\mathsf{C}_{U}(\mathbf{z}^{*})=\min_{\mathbf{z}\in\mathbb{R}^{n}}\mathsf{C}_{U}(\mathbf{z})$, neutral consensus equilibrium is also socially optimal. $\bullet$
@@ 6 Oscillatory Behavior of Opinions
In [31], we studied a special case of the dynamics proposed in the current paper. We showed through numerical simulations that in certain scenarios, the opinions exhibit oscillatory behavior. In this section, we analyze such behavior in the case of a pair of agents $\mathcal{V}=\{1,2\}$. For a general social network with $n$ agents, the analysis is significantly more complicated due to the higher dimension of the state space but also due to the complexity added by the social network. Thus, the analysis in the general case of a social network with $n$ agents would have to be a separate research work in itself and is out of the scope of this paper. Even though our analysis is restricted to two agents, it is still relevant since it can help us understand opinion behaviors or decision making in important systems such as two-party politics, duopoly economic markets etc.
@@
$\dot{z}_{1}=S_{1}(z_{1})+a_{12}\big{[}z_{2}-z_{1}\big{]},\ \dot{z}_{2}=S_{2}(z_{2})-a_{21}\big{[}z_{2}-z_{1}\big{]}.$ (17)
For the two-agent dynamics (17), it can be easily verified that if the opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit oscillatory behavior then they will have the same fundamental period of oscillations. We state this in the following result (proof in Appendix).
@@
(Equal period of oscillations) Consider the two-agent dynamics (17). Suppose the opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit oscillations with fundamental periods $T_{1}>0$ and $T_{2}>0$, respectively. Then, $T_{1}=T_{2}$. $\bullet$
Note that periodic orbits can exist for (17) only if at least one of the agents has an antagonist influence on the other as otherwise Theorem 4.3 guarantees convergence of opinions to the unique equilibrium from any arbitrary initial condition. In the next result (proof in Appendix), we give a necessary condition for the existence of periodic orbits for (17).
@@
(Necessary condition for existence of periodic solutions) Consider the two-agent dynamics (17). Opinions $z_{1}(t)$ and $z_{2}(t)$ exhibit periodic behavior only if all the following hold: (i) $a_{12}<0$ or $a_{21}<0$, (ii) $a_{12}a_{21}\neq 0$, and (iii) $(a_{21}+w_{1}+w_{2}+a_{12})<0$. $\bullet$
@@
(Geometric and other interpretations of the conditions for the existence of periodic opinions.) The condition given in Lemma 6.2 ensures existence of the following ellipse in the phase plane,
$\frac{z_{1}^{2}}{r_{1}}+\frac{z_{2}^{2}}{r_{2}}=-\frac{(w_{1}+w_{2}+a_{12}+a_{21})}{3}=:\upsilon>0.$ (18)
@@
The oscillatory behavior of opinions appears only when at least one of the agents has an antagonist influence on the other. This is similar to the so-called boomerang effect [34], where at least one of the agents has an antagonist influence on another agent. As a result, the agent on whom there is an antagonist influence shifts its opinion away from the other agent. In such situations, the opinions of the agents could converge to a disagreement equilibrium or could possibly keep on oscillating forever never converging to an equilibrium. $\bullet$
Now that we have dealt with necessary conditions for a periodic solution of (17) to exist, we conclude this section by giving sufficient conditions for the agents to exhibit periodic opinion profiles. More specifically, we give sufficient conditions on the conformity weights $a_{12}$, $a_{21}$ for a Hopf bifurcation to exist (proof in Appendix).
@@
(Existence of Hopf bifurcation) For the two-agent dynamics (17) with $a_{21}\in$ as the bifurcation parameter, let $\kappa_{i}:=\left[w_{i}+\frac{3m_{i}^{2}}{r_{i}}\right]>0,i\in\{1,2\}$. Let $m_{1}=m_{2}=m\in$ and $\left[a_{12}(\kappa_{2}-\kappa_{1})-\kappa_{1}^{2}\right]>0$. Then, a family of periodic orbits of (17) bifurcates from the consensus equilibrium $\mathbf{z}^{*}=m\mathbf{1}$ at $a_{21}=a_{21}^{*}:=-(\kappa_{1}+\kappa_{2}+a_{12})$. $\bullet$